RC Meeting and Runplan, Tuesday, Sept 25
- Beam last night was big at the target, and a Moller quad was found to have been turned on last evening during beam tuning. Latest scans taken this morning: https://logbooks.jlab.org/entry/3597363 https://logbooks.jlab.org/entry/3597364 https://logbooks.jlab.org/entry/3597276
- Restarted the big BPM calibration this morning: https://logbooks.jlab.org/entry/3597354
- Accelerator worked on getting high current to Hall C today. Last I heard, we can expect something around 15 uA (max stable) for this evening.
- Shifts will be manned normally starting OWL this evening until further notice.
1. Get beam into Hall with reasonable beam spot size, big BPM calibration, ion chamber calibration. 2. SHMS deflection test (approximately 2.5-6 hours, 10-20 uA). Call experts for step 4 when ramping Q3. 3. Check beam spot with harps, beam on yag with and without raster, carbon hole (5 uA) 4. Detector and trigger checkout (see run plan and detailed slides here: ). Calibration runs request 50 uA, 65 uA, and 20 uA (we will be capped tonight).
- Dave Meekins needs short access tomorrow morning to begin condensing the target (probably around 8am-ish)
- Jerry needs something from the hall and will coordinate with radcon
- Steve Lassiter is also going to take a look at the SHMS encoder
- Note: Hall A is planning a pass change for Thursday- there will be a few hours of no beam during that time.
- Encoder angles in the rotation GUI drift, and users can no longer modify the offsets themselves: https://logbooks.jlab.org/entry/3597240 https://logbooks.jlab.org/entry/3597512 https://logbooks.jlab.org/entry/3597362
- Epics IOCs moved back to cdaql1: https://logbooks.jlab.org/entry/3597491
- JP will do target training 10am Wednesday and 4pm Thursday. Contact him if interested!!